OpenStudy (anonymous):

I have to describe in paragraph how to simplify each expression 16 1/2 and the sqaure root of 100.......someone please help me!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

Well what is the square root of 100?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

No. \[100 = 50*2 \] \[50^2 = 2500\]

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Right. So since 10 squared is 100, the square root of 100 is 10. I'm not sure how you're expected to write a paragraph though.

OpenStudy (anonymous):

ok so how do you simplify 16 1/2

OpenStudy (anonymous):

is that 16 and 1/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, 16 is 32/2, so 16 + 1/2 = 32/2 + 1/2 = 33/2

OpenStudy (anonymous):

But again, I'm not sure what is supposed to take up the rest of the paragraph.

OpenStudy (anonymous):

ok thank you